# Dielectric Tensor Prediction for Inorganic Materials Using Latent Information from Preferred Potential

### v1

@@ Abstract
Dielectrics are materials with widespread applications in flash memory, central processing units, photovoltaics, capacitors, etc. However, the availability of public dielectric data remains limited, hindering research and development efforts. Previously, machine learning models focused on predicting dielectric constants as scalars, overlooking the importance of dielectric tensors in understanding material properties under directional electric fields for material design and simulation. This study demonstrates the value of common equivariant structural embedding features derived from a universal neural network potential in enhancing the prediction of dielectric properties. To integrate channel information from various-rank latent features while preserving the desired SE(3) equivariance to the second-rank dielectric tensors, we design an equivariant readout decoder to predict the total, electronic, and ionic dielectric tensors individually, and compare our model with the state-of-the-art models. Finally, we evaluate our model by conducting virtual screening on thermodynamical stable structure candidates in Materials Project. The material Ba2SmTaO6 with large band gaps ($E_{g}=3.36\mathrm{eV}$) and dielectric constants ($\epsilon=93.81$) is successfully identified out of the 14k candidate set. The results show that our methods give good accuracy on predicting dielectric tensors of inorganic materials, emphasizing their potential in contributing to the discovery of novel dielectrics.
Keywords SE(3), Graph neural networks, Inorganic materials, Dielectric tensor
@@ 1 Introduction
Dielectrics are an essential group of materials identified by their distinct electronic characteristics. They find extensive applications in various fields such as photovoltaics[1], energy storage[2, 3], microwave communication[4, 5], etc., in contemporary society. When subjected to an external electric field, a dielectric material is polarized by internally generating electric dipole moments. The dielectric constant $\epsilon$, also known as permittivity, is a measure of such dielectric effect described by the proportionality between the externally applied electric field to the generated field within the material:
$E_{i}^{\textrm{int}}=\sum_{j}\epsilon_{ij}^{-1}E_{j}^{\textrm{ext}}$ (1)
@@
where $E^{\textrm{int}}$ and $E^{\textrm{ext}}$ are the internal and external electric, respectively, and the indices $i,j\in\{1,2,3\}$ refer to the directions in the 3D space. The $3\times 3$ dielectric tensor has a minimum of 1 and a maximum of 6 independent elements for various types of systems due to the crystal symmetry. It can be represented by a sum of two components: the electronic contribution ($\epsilon^{\infty}$) and ionic contribution ($\epsilon^{0}$), which are generated by electron cloud distortion and atomic displacement, respectively, i.e., $\epsilon_{ij}=\epsilon^{\infty}_{ij}+\epsilon^{0}_{ij}$. Both high and low dielectric constants play an essential role in material design. For instance, high-$\kappa$ dielectrics enable higher-density charge storage to shrink device sizes, while low-$\kappa$ dielectrics are important to device packaging by minimizing the signal interference between neighboring interconnects.
@@
As of the present, only a limited number of materials in few hundreds have had their dielectric constants measured[6] because of the requirement of huge experimental efforts. In the past decades, first-principle calculations played a critical role for accelerating the discovery of new materials[7]. Density functional theory (DFT)[8, 9] has gained significant popularity due to its commendable balance between computational speed and accuracy for simulating materials in silico. This has, in turn, facilitated the emergence of a wide range of diverse new materials.[10, 11]. Similarly, density functional perturbation theory (DFPT) serves as a simulation tool for establishing the relationship between structure and dielectric properties, and the calculated data in the order of thousands are publicly accessible for material exploration[12, 6]. However, the DFT-based methods entail significant computational cost, making them less feasible for high-throughput screening, and it is only applicable to small systems with the number of atoms typically less than 1000[13], since its complexity scales poorly with the number of particles. Alternatively, machine learning is becoming an important research tool, providing an alternative method for quantitative structure?property relationship (QSPR) analysis for material science, There are numerous successful works in designing materials for various applications, including fluorescent molecules[14], electrets[15, 16], DDR1 kinase prohibitors[17], thermal-electric materials[18], etc.
@@
Several existing works made the efforts on making predictions for dielectric property directly from crystal/molecule descriptors by using the machine learning model. Umeda et al. adopted a random forest (RF) model to estimate experimental dielectric constants with the error range as large as 50% in the logarithmic scale on the purpose of developing ferroelectric material[19]. Takahashi et al. studied the prediction for static dielectric constants of metal oxides and analyzed importance of descriptors using RF models. Lin et al. employed a simple Gradient boosting regression (GBR) model to predict the electronic contribution of polycrystalline dielectric constants with RMSE of 0.824 and did analysis on feature importance for $\textrm{ABO}_{3}$-type compounds[20]. Similarly, Kim et al. used a GBR model to predict both the electronic and ionic contribution of $\textrm{ABO}_{3}$-type perovskites, achieving RMSE of 0.12 and 0.26 respectively[21]. Although these works proved the strong ability of machine learning models to rapidly estimate dielectric properties on unseen materials and are undoubtedly helpful to guide material designers before conducting wet experiments, there are still several limitations that hinder their full potential for exploration in the dielectric material space. Firstly, these tasks have mainly utilized simple machine learning models and manually selected descriptors for the models, thus unable to comprehensively showcase the structural characteristics of the materials which are crucial for more accurate dielectric tensor prediction. Additionally, these works only consider the specific types of inorganic materials (e.g. ABO3 perovskite structures) or materials containing specific combinations of elements. This narrow scope restricts the generalizability of the models and hinders their applicability to a wider range of dielectric materials. Lastly, they use a simple average scalar to express the dielectric constant by
$\epsilon=\sum_{i=1}^{3}\lambda_{i}/3$ (2)
@@
where $\lambda_{i}$ is the eigenvalue of $\boldsymbol{\epsilon}$. However, the off-diagonal elements in the dielectric tensor may reach large values and sometimes cannot be neglected in practical[22], thus prediction of the dielectric constants as tensors assists human researchers to understand the behavior of materials in the presence of a directed electric field. To distinguish, dielectric constants as scalars and as tensors are denoted with $\epsilon$ and $\boldsymbol{\epsilon}$, respectively. $\boldsymbol{\epsilon}$ could provide crucial comprehensive measures for designs of electronic devices in various applications.
@@
In recent years, graph neural networks (GNN) has been becoming a powerful tool for rapid atomic simulation in material science[23, 24] as graphs are a natural representation for crystals and molecules where nodes are atoms and edges are bonds between atom pairs. Many great GNN-based models[25, 26, 27, 28, 29, 30] are proposed and demonstrated their efficacy in accurately establishing a quantitative structure-property relationship. Recent works proposed the application of equivariant graph neural networks (EGNN) in materials science[31, 30] and discussed the necessity of equivariant representations to enhance model expressivity. TeaNet[31] is an EGNN architecture that satisfies special Euclidean group E(3) equivariance including translation, rotation and reflection. Based on TeaNet with certain modifications, PreFerred Potential (PFP)[32] is a universal neural network potential (NNP) model trained by massive data and has been released for multiple versions. What especially makes PFP robust for universal interatomic interactions is its enhanced representations of the local electronic environments centered around atoms and bonds by converting raw input data (atomic numbers and positions) to rank-0,1,2 tensors (i.e., scalars, vectors and tensors). Pretrained on the 22 million dataset (calculated by 1144 GPU years), the up-to-date PFP covers 72 elements in the periodic table for empirical universal interatomic potential simulation with strong transferability[33]. Its abundant intermediate interatomic information in pretrained layers of graph convolutional networks (GCN) and inclusion of equivariant rank-2 tensor features make it potentially suitable to transfer to diverse tensorial regression tasks with relatively limited datasets such as dielectric tensors.
@@
In this study, we propose Dielectric Tensor Neural Network(DTNN) to predict the three types of dielectric tensors, $\epsilon_{ij}^{\infty}$, $\epsilon_{ij}^{0}$ and $\epsilon_{ij}$, for inorganic materials across 72 supported elements with the equivairance to the input structure. By leveraging a pretrained PFP model as an efficient and expressive encoder, we demonstrated that the pretrained PFP processes a rich set of high-order latent compositional and structural information for tensorial property prediction. The ablation study shows that both latent atomic features and bond features from PFP contribute to the final prediction. To evaluate the performance of DTNN, we compared it against existing methods, PaiNN[30], M3GNet[34] and MatTen[35], on the dataset obtained from Materials Project (MP)[36]. Finally, we employed our model to perform a virtual screening on the design of dielectric materials for the microelectronic manufacturing application.
@@ 2.1 Overview
Figure 1: Diagram of the equivariant model for dielectric tensor prediction. The input graph is represented by initialized atom attributes $\mathcal{V}=\{(\mathbf{a}_{s},\mathbf{a}_{v},\mathbf{a}_{t})_{i}\}_{i=1:N^{a}}^{0}$ and bond attributes =$\mathcal{E}=\{(\mathbf{b}_{s_{\{i,j\}}},\mathbf{b}_{v_{\{i,j\}}})_{k}\}_{k=1:N^{b}}^{0}$ between atoms with indices of $i$ and $j$, where $N^{a}$ and $N^{b}$ are number of atoms and bonds. The intermediate knowledge output $\{(\mathbf{a}_{s},\mathbf{a}_{v},\mathbf{a}_{t})_{i}\}_{i=1:N^{a}}^{n}$ and $\{(\mathbf{b}_{s_{\{i,j\}}},\mathbf{b}_{v_{\{i,j\}}})_{k}\}_{k=1:N^{b}}^{n}$ from the $n$-th layer GCN of TeaNet/PFP is fed into an equivariant readout block to perform message interactions among different rank representations for a downstream task, i.e., dielectric tensor prediction here.
The challenge of GNN in materials science is that they usually require a large number of data to prevent models from overfitting, while such chemical data are generally expensive to collect either from simulation or experiments. Additionally, the numbers of calculated data for diverse properties can have large difference. For example, by the time of this work, there are more than $\sim$154k DFT-relaxed material structures with energies in MP, but only $\sim$7k (<5%) of them are available with dielectric constant data. In this work, we took a PFP as the parent encoder model, i.e., $N=5$. And for convenience, we denote the PFP with the $n(1\leq n\leq 5)$-th GCN layer as featurizer as PFP-L1, …,PFP-L5. Our method leverages a PFP model consisting of 5 GCN layers, which was pre-trained on 22 million first-principle calculations of potential energies, as a multi-order feature encoder to generate universal compositional and structural information for the second-order dielectric property prediction (see Figure 1). We design a separate model (see Fig.2) to capture interactions among the intermediate knowledge from PFP-$\mathrm{L}_{n}$ and make predictions on downstream tasks without the need to finetune parameters in the parent PFP model. This design ensures the effective utilization of high-order common structural information, leading to improved performance even with a relatively small region of available training data. Details about the module implementation are available in the Methods Section.
Figure 2: Schematic architecture of the equivariant dielectric model. (a) Overview of the DTNN framework; (b) Implementation of an equivariant block in DTNN; (c) Implementation of the readout block for dielectric tensor output.
@@ 2.2 Equivariance
Atomic systems are usually represented with coordinates in 3D Euclidean space. The same system could have transformations under the 3D symmetries: rotations, translations, and inversion. Under rotation in 3D space, certain properties of the system remain invariant, e.g., energy, meaning they do not change regardless of the rotation. Equivariant properties, including forces, stresses, and dielectric tensors, are expected to rotate accordingly with the system. In other words, as the system rotates, these properties undergo corresponding transformations to maintain their relationship with the system’s new orientation. Formally, we consider a function $\phi:X\rightarrow Y$ and a group transformation $G_{T}$ that acts on the vectors spaces $X$ and $Y$. The function $\phi$ is equivariant to the transformation $G_{T}$ if the following condition holds for every element $x\in X$: $\phi(G_{T}(x))=G_{T}^{\prime}(\phi(x))$, where $G_{T}(x)$ represents the group transformation applied to the input $x$, and $G_{T}^{\prime}(f(x))$ represents a related transformation applied to the output $\phi(x)$. Here, we focus on the rotational equivariance of the second-rank dielectric tensor $\boldsymbol{\epsilon}$ with respect to the input vector space $\mathcal{X}$, (Fig.3), whose rotational equivariant condition is $R\Phi(\mathcal{X})R^{T}=\Phi(R\mathcal{X})$ where $R$ is the arbitrary orthogonal rotation matrix and $\Phi$ is the neural networks to map $\mathcal{X}$ to $\boldsymbol{\epsilon}$.
Figure 3: Rotational equivariance on dielectric constants of inorganic materials. $\overrightarrow{E}$ is the external electric field, $\overrightarrow{D}$ is the corresponding electric displacement, and $\boldsymbol{\epsilon}$ is the $3\times 3$ dielectric tensor.

### v2

@@ Abstract
Dielectrics are crucial for technologies like flash memory, CPUs, photovoltaics, and capacitors, but public data on these materials are scarce, restricting research and development. Existing machine learning models have focused on predicting scalar polycrystalline dielectric constants, neglecting the directional nature of dielectric tensors essential for material design. This study leverages multi-rank equivariant structural embeddings from a universal neural network potential to enhance predictions of dielectric tensors. We develop an equivariant readout decoder to predict total, electronic, and ionic dielectric tensors while preserving O(3) equivariance, and benchmark its performance against state-of-the-art algorithms. Virtual screening of thermodynamically stable materials from Materials Project for two discovery tasks, high-dielectric and highly anisotropic materials, identifies promising candidates including Cs2Ti(WO4)3 (band gap $E_{g}=2.93\mathrm{eV}$, dielectric constant $\varepsilon=180.90$) and CsZrCuSe3 (anisotropic ratio $\alpha_{r}=121.89$). The results demonstrate our model’s accuracy in predicting dielectric tensors and its potential for discovering novel dielectric materials.
Keywords O(3) equivariance, Graph neural networks, Inorganic materials, Dielectric tensor
@@ 1 Introduction
Dielectrics are an essential group of materials identified by their distinct electronic characteristics. They find extensive applications in various fields such as photovoltaics [1], energy storage [2, 3], microwave communication [4, 5], etc., in contemporary society. When subjected to an external electric field, a dielectric material is polarized by internally generating electric dipole moments. The dielectric constant $\varepsilon$, also known as permittivity, is a measure of such dielectric effect described by the proportionality between the externally applied electric field to the generated field within the material:
$E_{i}^{\textrm{int}}=\sum_{j}\varepsilon_{ij}^{-1}E_{j}^{\textrm{ext}}$ (1)
@@
where $E^{\textrm{int}}$ and $E^{\textrm{ext}}$ are the internal and external electric, respectively, and the indices $i,j\in\{1,2,3\}$ refer to the directions in the 3D space. The $3\times 3$ dielectric tensor has a minimum of 1 and a maximum of 6 independent elements for various types of systems due to the crystal symmetry. It can be represented by a sum of two components: the electronic contribution ($\varepsilon^{\infty}$) and ionic contribution ($\varepsilon^{0}$), which are generated by electron cloud distortion and atomic displacement, respectively, i.e., $\varepsilon_{ij}=\varepsilon^{\infty}_{ij}+\varepsilon^{0}_{ij}$. Both high and low dielectric constants play an essential role in material design. For instance, high-$\kappa$ dielectrics enable higher-density charge storage to shrink device sizes, while low-$\kappa$ dielectrics are important to device packaging by minimizing signal interference between neighboring interconnects.
@@
As of the present, only a limited number of materials in few hundreds have had their dielectric constants measured [6] because of the requirement of huge experimental efforts. In the past decades, first-principle calculations played a critical role for accelerating the discovery of new materials [7]. Density functional theory (DFT) [8, 9] has gained significant popularity due to its commendable balance between computational speed and accuracy for simulating materials in silico. This has, in turn, facilitated the emergence of a wide range of diverse new materials [10, 11]. Similarly, density functional perturbation theory (DFPT) serves as a simulation tool for establishing the relationship between structure and dielectric properties, and the calculated data in the order of thousands are publicly accessible for material exploration [12, 6]. However, the DFT-based methods entail significant computational cost, making them less feasible for high-throughput screening, and it is only applicable to small systems with the number of atoms typically less than 1000 [13], since its complexity scales poorly with the number of particles. Alternatively, machine learning is becoming an important research tool, providing an alternative method for quantitative structure property relationship (QSPR) analysis for material science, There are numerous successful works in designing materials for various applications, including fluorescent molecules [14], electrets [15, 16], DDR1 kinase prohibitors [17], thermal-electric materials [18], etc.
@@
Several existing works made efforts to make predictions for dielectric property directly from crystal/molecule descriptors by using the machine learning model. Umeda et al. adopted a random forest (RF) model to estimate experimental dielectric constants with an error range as large as 50% in the logarithmic scale on the purpose of developing ferroelectric material [19]. Takahashi et al. studied the prediction for static dielectric constants of metal oxides and analyzed the importance of descriptors using RF models. Lin et al. employed a simple Gradient boosting regression (GBR) model to predict the electronic contribution of polycrystalline dielectric constants with RMSE of 0.824 and did analysis on feature importance for $\textrm{ABO}_{3}$-type compounds [20]. Similarly, Kim et al. used a GBR model to predict both the electronic and ionic contribution of $\textrm{ABO}_{3}$-type perovskites, achieving RMSE of 0.12 and 0.26 respectively [21]. Although these works proved the strong ability of machine learning models to rapidly estimate dielectric properties on unseen materials and are undoubtedly helpful to guide material designers before conducting wet experiments, there are still several limitations that hinder their full potential for exploration in the dielectric material space. These tasks focus on the average polycrystalline dielectric constants approximated from the dielectric tensor, given by
$\varepsilon=\sum_{i=1}^{3}\lambda_{i}/3$ (2)
@@
where $\lambda_{i}$ is the eigenvalue of $\boldsymbol{\varepsilon}$. The Matbench dataset [22] includes the dielectric task for benchmarking various material property models. The property in this task is the refractive index ($\eta$) which is related to the electronic dielectric constant, as given by $\eta=\sqrt{\varepsilon^{\infty}}$. Various GNN-based models are evaluated on this task. However, these works ignored the inherent anisotropy in crystal materials. Additionally, the off-diagonal elements in the dielectric tensor may reach large values and sometimes cannot be neglected in practical [23], thus prediction of the dielectric constants as tensors assists human researchers to understand the behavior of materials in the presence of a directed electric field. To distinguish, dielectric constants as scalars and as tensors are denoted with $\varepsilon$ and $\boldsymbol{\varepsilon}$, respectively. $\boldsymbol{\varepsilon}$ could provide crucial comprehensive measures for the designs of electronic devices in various applications.
@@
In recent years, graph neural networks (GNN) has been becoming a powerful tool for rapid atomic simulation in material science [24, 25] as graphs are a natural representation for crystals and molecules where nodes are atoms and edges are bonds between atom pairs. Many great GNN-based models [26, 27, 28, 29, 30, 31] are proposed and demonstrated their efficacy in accurately establishing a quantitative structure-property relationship. Recent works proposed the application of equivariant graph neural networks (EGNN) in materials science [32, 31] and discussed the necessity of equivariant representations to enhance model expressivity. TeaNet [32] is an EGNN architecture that satisfies special Euclidean group E(3) equivariance including translation, rotation and reflection. Based on TeaNet with certain modifications, PreFerred Potential (PFP[33] is a universal neural network potential (NNP) model trained by massive data and has been released for multiple versions. What especially makes PFP robust for universal interatomic interactions is its enhanced representations of the local electronic environments centered around atoms and bonds by converting raw input data (atomic numbers and positions) to rank-0,1,2 tensors (i.e., scalars, vectors and tensors). Pretrained on the original dataset with 22 million diverse structures (calculated by 1144 GPU years), the up-to-date PFP covers 72 elements in the periodic table for empirical universal interatomic potential simulation with strong transferability [34]. We refer readers to [33] for more details about this original training dataset across all systems including molecular, crystal, slab, cluster, adsorption, and disordered structures. Its abundant intermediate interatomic information in pretrained layers of graph convolutional networks (GCN) and inclusion of equivariant rank-2 tensor features make it potentially suitable to transfer to diverse tensorial regression tasks with relatively limited datasets such as dielectric tensors. During the same period as this study, Lou et al. [35] proposed using equivariant graph neural networks to learn irreducible representations for predicting dielectric materials with highly anisotropic electronic dielectric tensors. In contrast, our study leverages transfer learning of pretrained geometric features across properties with varying tensor orders to enable dielectric tensor predictions over a broader numerical range. Moreover, instead of focusing exclusively on electronic components, we evaluate our models on electronic, ionic, and total dielectric tasks, demonstrating state-of-the-art performance in predicting both scalar and tensor properties.
@@
In this study, we propose Dielectric Tensor Neural Network(DTNet) to predict the three types of dielectric tensors, $\varepsilon_{ij}^{\infty}$, $\varepsilon_{ij}^{0}$ and $\varepsilon_{ij}$, for inorganic materials across 72 supported elements with the equivairance to the input structure. By leveraging a pretrained PFP model as an efficient and expressive encoder, we demonstrated that the pretrained PFP processes a rich set of high-order latent compositional and structural information for tensorial property prediction. The ablation study shows that both latent atomic features and bond features from PFP contribute to the final prediction. To evaluate the performance of DTNet, we compared it against existing methods, PaiNN [31], M3GNet [36] and MatTen [37], on the dataset obtained from Materials Project (MP) [38]. Finally, we applied our model to virtual screening for two discovery tasks: identifying high-dielectric materials and highly anisotropic dielectrics.
@@ 2.1 Overview
Figure 1: Diagram of the equivariant model for dielectric tensor prediction. The input graph is represented by initialized atom attributes $\mathcal{V}=\{(\boldsymbol{a}_{s},\boldsymbol{a}_{v},\boldsymbol{a}_{t})_{i}\}_{i=1:N^{a}}^{0}$ and bond attributes =$\mathcal{E}=\{(\boldsymbol{b}_{s,{\{i,j\}}},\boldsymbol{b}_{v,{\{i,j\}}})_{k}\}_{k=1:N^{b}}^{0}$ between atoms with indices of $i$ and $j$, where $N^{a}$ and $N^{b}$ are number of atoms and bonds. The intermediate knowledge output $\{(\boldsymbol{a}_{s},\boldsymbol{a}_{v},\boldsymbol{a}_{t})_{i}\}_{i=1:N^{a}}^{n}$ and $\{(\boldsymbol{b}_{s,{\{i,j\}}},\boldsymbol{b}_{v,{\{i,j\}}})_{k}\}_{k=1:N^{b}}^{n}$ from the $n$-th layer GCN of PFP is fed into an equivariant readout block to perform message interactions among different rank representations for a downstream task, i.e., dielectric tensor prediction here.
The challenge of GNN in materials science is that they usually require a large number of data to prevent models from overfitting, while such chemical data are generally expensive to collect either from simulation or experiments. Additionally, the numbers of calculated data for diverse properties can have large difference. For example, by the time of this work, there are more than $\sim$154k DFT-relaxed material structures with energies in MP, but only $\sim$7k (<5%) of them are available with dielectric constant data. In this work, we took a PFP as the parent encoder model, i.e., $N=5$. And for convenience, we denote the PFP with the $n(1\leq n\leq 5)$-th GCN layer as featurizer as PFP-L1, …,PFP-L5. Our method leverages a PFP model consisting of 5 GCN layers, which was pre-trained on 22 million first-principle calculations of potential energies, as a multi-order feature encoder to generate universal compositional and structural information for the second-order dielectric property prediction (see Figure 1). We design a separate model (see Fig.2) to capture interactions among the intermediate knowledge from PFP-$\mathrm{L}_{n}$ and make predictions on downstream tasks without the need to finetune parameters in the parent PFP model. This design ensures the effective utilization of high-order common structural and compositional information, leading to improved performance even with a relatively small region of available training data. Details about the module implementation are available in the Methods Section.
Figure 2: Schematic architecture of the equivariant dielectric model DTNet. (a) Overview of the DTNet framework; (b) Implementation of an equivariant block in DTNet; (c) Implementation of the readout block for dielectric tensor output
@@ 2.2 Equivariance
Atomic systems are usually represented with coordinates in 3D Euclidean space. The same system could have transformations under the 3D symmetries: rotations, translations, and inversion. Under orthogonal transformations in 3D space, certain properties of the system remain invariant, e.g., energy, meaning they do not change regardless of the rotation. Equivariant properties, including forces, stresses, and dielectric tensors, are expected to transform accordingly with the system. In other words, as the system rotates or reflects, these properties undergo corresponding transformations to maintain their relationship with the system’s new orientation. Formally, we consider a function $\phi:X\rightarrow Y$ and a group transformation $G_{T}$ that acts on the vectors spaces $X$ and $Y$. The function $\phi$ is equivariant to the transformation $G_{T}$ if the following condition holds for every element $x\in X$: $\phi(G_{T}(x))=G_{T}^{\prime}(\phi(x))$, where $G_{T}(x)$ represents the group transformation applied to the input $x$, and $G_{T}^{\prime}(f(x))$ represents a related transformation applied to the output $\phi(x)$. Here, we focus on the rotation/inversion equivariance of the second-rank dielectric tensor $\boldsymbol{\varepsilon}$ with respect to the input vector space $\mathcal{X}$ (Fig.3), whose equivariant condition is $R\Phi(\mathcal{X})R^{T}=\Phi(R\mathcal{X})$ where $R$ is the arbitrary orthogonal matrix and $\Phi$ is the neural networks to map $\mathcal{X}$ to $\boldsymbol{\varepsilon}$.
Figure 3: Rotational equivariance on dielectric constants of inorganic materials. $\overrightarrow{E}$ is the external electric field, $\overrightarrow{D}$ is the corresponding electric displacement, and $\boldsymbol{\varepsilon}$ is the $3\times 3$ dielectric tensor.